Consider the following projects: Project A has a cost of 75 and uniform annual benefit of 18.8; Project B has a cost of 50 and uniform annual benefit of 13.9; Project C has a cost of 15 and uniform annual benefit of 4.5; Project D has a cost of 90 and uniform annual benefit of 23.8. Using a 5-year analysis with MARR = 10%, which project should be selected using the payback period.

Group of answer choices

Project A

Project D

Project B

Project C

7.

MARR stands for “Minimum Attractive Rate of Return”.

Group of answer choices

True

False
Answer #1

1.

B/C of A = 18.8 * (P/A,10%,5) / 75

= 18.8 * 3.790787 / 75

= 0.95

B/C of B = 13.9 * (P/A,10%,5) / 50

= 13.9 * 3.790787 / 50

= 1.05

B/C of C = 4.5 * (P/A,10%,5) / 15

= 4.5 * 3.790787 / 15

= 1.14

B/C of D = 23.8 * (P/A,10%,5) / 90

= 23.8 * 3.790787 / 90

= 1.002

Only B, C & D qualify for incremental analysis

B/C of (B-C) = (13.9 - 4.5)* (P/A,10%,5) / (50 - 15)

= 9.4 * 3.790787 / 35

= 1.02

B is selected and C is rejected

B/C of (D - B) = (23.8 - 13.9)* (P/A,10%,5) / (90-50)

= 9.9 * 3.790787 / 40

= 0.94

B is selected and D is rejected

Finally B should be selected as per incremental B/C analysis

7.

MARR = Minimum Attractive Rate of Return

Given statement is TRUE